A runner starts a race at the starting line. A second runner starts at 44 yards. If the first runner moves 7 yards each second and the second runner moves 3 yards each second, at what distance will the first runner catch the second runner?
Answer by lwsshak3(11628)

A runner starts a race at the starting line. A second runner starts at 44 yards. If the first runner moves 7 yards each second and the second runner moves 3 yards each second, at what distance will the first runner catch the second runner?
..
let x = distance the first runner traveled
then x-44 = distance the second runner traveled
time traveled = distance/speed
Both traveled the same amount of time
x/7=(x-44)/3
3x=7x-(7)(44)
-4x=-(7)(44)
x=77 yards
ans:
The first runner will have traveled a distance of 77 yards to catch up with the second runner